# Electrical conductivity of randomly placed linear wires: a mean field approach

### v2

@@ Abstract
Using the mean-field approximation, a formula for the effective electrical conductivity of a two-dimensional system of randomly arranged conducting sticks with a given orientation distribution was obtained. Both the resistance of the sticks themselves and the resistance of the contacts between them were taken into account. The accuracy in the resulting formula was analyzed. A comparison of the theoretical predictions of mean-field approach with the results of direct electrical conductivity calculations for several model orientation distributions describing systems with crossed sticks demonstrated good agreement. Our study showed that cross-alignment of nanowires should lead to a decreasing in the electrical conductivity compared to electrodes with isotropically arranged nanowires. We suppose that the widely used model with zero-width sticks is quite acceptable for systems of cross-aligned nanowires, but overestimates their connectivity in isotropic systems. Thus, the enhancement of the electrical conductivity of conducting films with cross-aligned nanowires may be due to a significant difference in the network topology.
@@ 1 Introduction
Modern technologies make it possible to produce conductive films consisting of two layers of nanowires: in the first layer, the nanowires are aligned predominantly along one direction, while in the second layer, they are aligned perpendicularly. Examples of such technologies are dip coating [1], the capillary printing technique [2], blade-coating method [3], bar-coating assembly [4], direct-writing solution process guided by a conical fiber array [5], deposition by using the shear force from the high-speed rotating concentrated nanowire solution [6] or from Mayer’s rod [7, 8, 9], a continuous roll-to-roll process [10]. The optoelectrical performance of films with a cross-aligned arrangement of metallic nanowires is superior to that of those with a random arrangement [11]. The cross-aligned nanowires are favorable for their nanowelding, which plays a crucial role in reducing junction resistance between nanowires in electrodes base on metallic nanowires [8, 9]. Conductive films based on cross-aligned nanowires are more uniform compared to films based on randomly arranged nanowires, which reduces the likelihood of electrical breakdown and the appearance of hot spots (overheated areas) [4]. The surface roughness of these films is less as compared to films based on randomly arranged nanowires [1, 4]. Experiments show that for the same transparency, which is determined by the concentration of nanowires, systems with cross-aligned nanowires have lower sheet resistance compared to isotropic systems [3, 4].
@@ 2.1 Model
Let $N$ identical zero-width conductive sticks of length $l$ be placed within a square domain $\mathcal{D}$ of size $L\times L$ ($L>l$) with periodic boundary conditions (PBCs). The number density of sticks, i.e., the number of sticks per unite area, is
@@ 2.3 Simulation
Conductive sticks were mimicked by linear segments of length $l=1$. These segments were placed within a square domain of linear size $L=32$. Periodic boundary conditions were applied to reduce the finite size effect. The centres of the segments were randomly placed within the domain. Their orientations of segments obeyed to one of the desired ODFs. In such a way, a network was obtained, which served as a model of conducting films based on cross-aligned nanowires.
@@
We performed test calculations to determine the finite size effect and effect of the value of the parameter $\Delta$. For values of $\Delta=10^{-6}$ and $\Delta=10^{-3}$ ($L=32$), the electrical conductivities were consistent within the statistical error for all values of the number density (Fig. 2). In such a way, even $\Delta=10^{-3}$ can be considered as a limiting case $\Delta\ll 1$. By contrast, despite the PBCs, the finite size effect is fairly visible, namely, the electrical conductivity increases as the system sizes enlarges.
@@ 3.1 Analytical solution for the effective conductivity
For convenience of further analysis, the effective electrical conductivity is written in a dimensionless form. Here, $N_{j}$ is the number of intersection of sticks with the given stick, $f(N_{\text{j}},N,p)$ is the fraction of sticks with exactly $N_{\text{j}}$ contacts. The number of intersections per stick is supposed to obey the binomial distribution $N_{\text{j}}\sim\operatorname{Bin}(N,p)$,
@@
$\lambda_{1}\propto 1+\sqrt{\frac{\Delta}{N_{\text{j}}}},$
$A=\frac{\left(\lambda_{1}-1\right)\left(\lambda_{1}^{{N_{\text{j}}}+1}-1\right)}{\lambda_{1}\left(\lambda_{1}^{N_{\text{j}}}+1\right)}\propto\lambda_{1}-1\propto\sqrt{\frac{\Delta}{N_{\text{j}}}}.$
@@
$\mathbb{E}\left[A\right]\approx\sqrt{\frac{\Delta}{\langle N_{\text{j}}\rangle}}\left(1+\frac{3}{8\langle N_{\text{j}}\rangle}\right),$
@@
$\sqrt{\frac{\Delta}{\langle N_{\text{j}}\rangle}}.$
@@
Figure 3 demonstrates that the across-alignment of sticks slightly decreases the backbone fraction.
@@ 3.2 Comparison of the analytical solution with simulations
Figure 5 compares our the direct computations of the electrical conductivity for the irregular cross with the predictions of the MFA using the ODF (5). Two values of the parameter $\varepsilon$ were used to produce irregular cross-aligned systems of nanowires. Thus, $\varepsilon=\pi/32$ corresponds to a narrow distribution, while $\varepsilon=\pi/8$ gets a wider distribution of nanowire orientations (see Fig. 1). The electrical conductivity computed in Ref. [11] is also presented for comparison. Since the ODF (3) was used in Ref. [11], this can only be a comparison of trends, not a quantitative comparison. Figure 5 suggests that the width of the distribution has a small effect on the electrical conductivity, especially in the cases when the junction resistance is not a dominating factor. When the junction resistance is a dominating factor, the wider distribution the smaller the electrical conductivity.
@@
Figure 7 suggests that the asymptotic behavior of the electrical conductivity is insensitive to the particular ODF. The cross-alignment demonstrates no visible effect on the electrical conductivity when $\Delta\gg 1$. The effect of cross-alignment on the electrical conductivity is negligible when $\Delta\gg 1$. The electrical conductivity slightly decreases from isotropic distribution to the cross-alignment when $\Delta\ll 1$.
@@ 4 Conclusion
Using the mean-field approximation, we rigorously derived a formula for the effective electrical conductivity of a 2D system of randomly arranged conducting rods with a given orientation distribution. The formula takes into account both the resistance of the rods themselves and the resistance of the contacts between them. For practical application, the formula was simplified by replacing the mean value of the certain function with the function of the mean value of its argument. An accuracy estimate for this resulting formula demonstrated that such a replacement leads to a slight overestimate of the electrical conductivity. Theoretical predictions of the mean-field approximation were compared with the results of direct conductivity calculations for several model orientation distributions describing systems with crossed rods. The good agreement between the theoretical estimates and the direct calculation results demonstrates the feasibility of using the mean-field approximation to describe the electrical conductivity of conducting transparent films with crossed nanowires. Neither the calculations nor the mean-field theory predictions confirm the results of [11] for the case where junction resistance dominates. Electrical conductivity does not depend on the specific type of the orientational distribution, but only on the average modulus of the sine of the angle between the conductors. Cross-alignment reduces electrical conductivity when junction resistance dominates over the nanowire resistance and has no significant effect in other cases. Cross-alignment leads to a slight increase in the percolation threshold, which is negative for the electrical conductivity. Cross-alignment slightly reduces the fraction of conductors belonging to the backbone of the percolation cluster, thus decreases the electrical conductivity. Summarizing, all considered effects of the nanowire cross-alignment have to reduce the electrical conductivity of transparent conductive films as compared to the films where the orientation of nanowires is equiprobable. Since experiments show the opposite, one can be concluded that the real cause lies outside the model. One of the possible explanations was proposed in Ref. [11]. The authors explain the observed effect due to decrease of the junction resistance in the case of cross-aligned nanowires. However, this explanation is not the only possible one. Although the ratio of the nanowires’ length to their width is large, a zero-width sticks is an extremely simplified model. Using zero-width sticks results in the actual quasi-three-dimensional (Q3D) network of nanowires being modeled using a 2D network with completely different properties [12]. However, the Q3D model also represents a limiting case, which is unlikely to be realized in the real world. In this model, nanowires are assumed to be infinitely rigid (hard rods of finite width), meaning they are incapable of any deformation. By contrast, using zero-width sticks is somewhat equivalent to treating nanowires as infinitely soft, meaning they are capable of perfectly bending around each other (soft threads). Obviously, real nanowires exhibit some intermediate properties: they can be deformed, but this deformation does not ensure perfect bending. We suppose that both 2D and Q3D networks are irrelevant models for isotropic nanowire networks. In the case of an isotropic orientation distribution, the 2d model significantly overestimates the average number of contacts on a conductor compared to the Q3D model. This overestimation is particularly noticeable when junction resistance dominate over the wire resistance. However, the 2D network appears to be quite acceptable to model the two-layered transparent electrodes where all the nanowires in the first layer are aligned predominantly along one direction and in the perpendicular direction in the second layer.
@@ Appendix A Detailed derivation of the master equation
Let the ODF of the conductor segment lengths be $f_{l}(i_{k};N_{\text{j}})$. Multiply equation (30) by $f_{l}(i_{k};N_{\text{j}})$ and integrate from 0 to $l$, then

### v3

@@ Abstract
Using the mean-field approximation, a formula for the effective electrical conductivity of a two-dimensional system of randomly arranged conducting sticks with a given orientation distribution was obtained. Both the resistance of the sticks themselves and the resistance of the contacts between them were taken into account. The accuracy in the resulting formula was analyzed. A comparison of the theoretical predictions of mean-field approach with the results of direct electrical conductivity calculations for several model orientation distributions describing systems with crossed sticks demonstrated good agreement. Our study showed that cross-alignment of nanowires should lead to a decrease in the electrical conductivity compared to electrodes with isotropically arranged nanowires. We suppose that the widely used model with zero-width sticks is quite acceptable for systems of cross-aligned nanowires, but overestimates their connectivity in isotropic systems. Thus, the enhancement of the electrical conductivity of conducting films with cross-aligned nanowires may be due to a significant difference in the network topology.
@@ 1 Introduction
Modern technologies make it possible to produce conductive films consisting of two layers of nanowires: in the first layer, the nanowires are aligned predominantly along one direction, while in the second layer, they are aligned perpendicularly. Examples of such technologies are dip coating [1], the capillary printing technique [2], blade-coating method [3], bar-coating assembly [4], direct-writing solution process guided by a conical fiber array [5], deposition by using the shear force from the high-speed rotating concentrated nanowire solution [6] or from Mayer’s rod [7, 8, 9], a continuous roll-to-roll process [10]. The optoelectrical performance of films with a cross-aligned arrangement of metallic nanowires is superior to that of those with a random arrangement [11]. The cross-aligned nanowires are favorable for their nanowelding, which plays a crucial role in reducing junction resistance between nanowires in electrodes based on metallic nanowires [8, 9]. Conductive films based on cross-aligned nanowires are more uniform compared to films based on randomly arranged nanowires, which reduces the likelihood of electrical breakdown and the appearance of hot spots (overheated areas) [4]. The surface roughness of these films is less as compared to films based on randomly arranged nanowires [1, 4]. Experiments show that for the same transparency, which is determined by the concentration of nanowires, systems with cross-aligned nanowires have lower sheet resistance compared to isotropic systems [3, 4].
@@ 2.1 Model
Let $N$ identical zero-width conductive sticks of length $l$ be placed within a square domain $\mathcal{D}$ of size $L\times L$ ($L>l$) with periodic boundary conditions (PBCs). The number density of sticks, i.e., the number of sticks per unit area, is
@@ 2.3 Simulation
Conductive sticks were mimicked by linear segments of length $l=1$. These segments were placed within a square domain of linear size $L=32$. Periodic boundary conditions were applied to reduce the finite size effect. The centres of the segments were randomly placed within the domain. Their orientations of segments obeyed one of the desired ODFs. In such a way, a network was obtained, which served as a model of conducting films based on cross-aligned nanowires.
@@
We performed test calculations to determine the finite size effect and effect of the value of the parameter $\Delta$. For values of $\Delta=10^{-6}$ and $\Delta=10^{-3}$ ($L=32$), the electrical conductivities were consistent within the statistical error for all values of the number density (Fig. 2). In such a way, even $\Delta=10^{-3}$ can be considered as a limiting case $\Delta\ll 1$. By contrast, despite the PBCs, the finite size effect is fairly visible, namely, the electrical conductivity increases as the system size enlarges.
@@ 3.1 Analytical solution for the effective conductivity
For convenience of further analysis, the effective electrical conductivity is written in a dimensionless form. Here, $N_{j}$ is the number of intersections of sticks with the given stick, $f(N_{\text{j}},N,p)$ is the fraction of sticks with exactly $N_{\text{j}}$ contacts. The number of intersections per stick is supposed to obey the binomial distribution $N_{\text{j}}\sim\operatorname{Bin}(N,p)$,
@@
$\lambda_{1}\propto 1+\sqrt{\frac{\Delta}{N_{\text{j}}}},$ (17)
$A=\frac{\left(\lambda_{1}-1\right)\left(\lambda_{1}^{{N_{\text{j}}}+1}-1\right)}{\lambda_{1}\left(\lambda_{1}^{N_{\text{j}}}+1\right)}\propto\lambda_{1}-1\propto\sqrt{\frac{\Delta}{N_{\text{j}}}}.$ (18)
@@
$\mathbb{E}\left[A\right]\approx\sqrt{\frac{\Delta}{\langle N_{\text{j}}\rangle}}\left(1+\frac{3}{8\langle N_{\text{j}}\rangle}\right),$ (19)
@@
$\sqrt{\frac{\Delta}{\langle N_{\text{j}}\rangle}}.$ (20)
@@
Figure 3 demonstrates that the across-alignment of sticks slightly decreases the backbone fraction. Thus, at the same number density of nanowires, the fraction of conductors participating in electrical conductivity is higher in the case of a system with an equally probable orientation of nanowires than in the case of a system with cross-aligned nanowires. This means that, within the framework of a two-dimensional nanowire network model, orientational order leads to a decrease in electrical conductivity.
@@ 3.2 Comparison of the analytical solution with simulations
Figure 5 compares our direct computations of the electrical conductivity for the irregular cross with the predictions of the MFA using the ODF (5). Two values of the parameter $\varepsilon$ were used to produce irregular cross-aligned systems of nanowires. Thus, $\varepsilon=\pi/32$ corresponds to a narrow distribution, while $\varepsilon=\pi/8$ gets a wider distribution of nanowire orientations (see Fig. 1). The electrical conductivity computed in Ref. [11] is also presented for comparison. Since the ODF (3) was used in Ref. [11], this can only be a comparison of trends, not a quantitative comparison. Figure 5 suggests that the width of the distribution has a small effect on the electrical conductivity, especially in the cases when the junction resistance is not a dominating factor. When the junction resistance is a dominating factor, the wider distribution the smaller the electrical conductivity.
@@
Figures 4, 5, and 6 show that, in full agreement with our estimate of the error that arises when replacing the mean value of a function with a function of the mean value of its argument (18), (19), and (20), the MFA overestimates the value of electrical conductivity the more the smaller the value of the delta parameter.
Figure 7 suggests that the asymptotic behavior of the electrical conductivity is insensitive to the particular ODF. The cross-alignment demonstrates no visible effect on the electrical conductivity when $\Delta\gg 1$. The effect of cross-alignment on the electrical conductivity is negligible when $\Delta=1$. The electrical conductivity slightly decreases from isotropic distribution to the cross-alignment when $\Delta\ll 1$.
@@ 4 Conclusion
Using the mean-field approximation, we rigorously derived a formula for the effective electrical conductivity of a 2D system of randomly arranged conducting rods with a given orientation distribution. The formula takes into account both the resistance of the rods themselves and the resistance of the contacts between them. For practical application, the formula was simplified by replacing the mean value of the certain function with the function of the mean value of its argument. An accuracy estimate for this resulting formula demonstrated that such a replacement leads to a slight overestimate of the electrical conductivity. Theoretical predictions of the mean-field approximation were compared with the results of direct conductivity calculations for several model orientation distributions describing systems with crossed rods. The good agreement between the theoretical estimates and the direct calculation results demonstrates the feasibility of using the mean-field approximation to describe the electrical conductivity of conducting transparent films with crossed nanowires. Neither the calculations nor the mean-field theory predictions confirm the results of [11] for the case where junction resistance dominates. Electrical conductivity does not depend on the specific type of the orientational distribution, but only on the average modulus of the sine of the angle between the conductors. Cross-alignment reduces electrical conductivity when junction resistance dominates over the nanowire resistance and has no significant effect in other cases. Cross-alignment leads to a slight increase in the percolation threshold, which is negative for the electrical conductivity. Cross-alignment slightly reduces the fraction of conductors belonging to the backbone of the percolation cluster, thus decreasing the electrical conductivity. Summarizing, all considered effects of the nanowire cross-alignment have to reduce the electrical conductivity of transparent conductive films as compared to the films where the orientation of nanowires is equiprobable. Since experiments show the opposite, one can be concluded that the real cause lies outside the model. One of the possible explanations was proposed in Ref. [11]. The authors explain the observed effect due to decrease of the junction resistance in the case of cross-aligned nanowires. However, this explanation is not the only possible one. Although the ratio of the nanowires’ length to their width is large, a zero-width stick is an extremely simplified model. Using zero-width sticks results in the actual quasi-three-dimensional (Q3D) network of nanowires being modeled using a 2D network with completely different properties [12]. It was shown [12] that in the 2D model, the average number of intersections per nanowire increases linearly with increasing number density of nanowires, whereas in the Q3D model, the average number of intersections is almost independent of the number density of nanowires. In real-world systems of randomly distributed and equally probable oriented nanowires, the average number of contacts is expected to grow with increasing number density of nanowires more slowly than linearly. Thus, the 2D model overestimates the connectivity of the nanowire network, leading to an overestimated electrical conductivity in both direct calculations and estimates within the MFA. On the other hand, in bilayer systems, such as those with cross-aligned nanowires, the 2D model is quite correct. However, the Q3D model also represents a limiting case, which is unlikely to be realized in the real world. In this model, nanowires are assumed to be infinitely rigid (hard rods of finite width), meaning they are incapable of any deformation. By contrast, using zero-width sticks is somewhat equivalent to treating nanowires as infinitely soft, meaning they are capable of perfectly bending around each other (soft threads). Obviously, real nanowires exhibit some intermediate properties: they can be deformed, but this deformation does not ensure perfect bending. We suppose that both 2D and Q3D networks are irrelevant models for isotropic nanowire networks. In the case of an isotropic orientation distribution, the 2D model significantly overestimates the average number of contacts on a conductor compared to the Q3D model. This overestimation is particularly noticeable when junction resistance dominates over the wire resistance. However, the 2D network appears to be quite acceptable to model the two-layered transparent electrodes where all the nanowires in the first layer are aligned predominantly along one direction and in the perpendicular direction in the second layer.
@@ Acknowledgements
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author contributions
YYT: Supervision (lead), Conceptualization (lead), Methodology (lead), Visualization (lead), Writing – Original Draft Preparation (lead), Writing – Review & Editing (lead); AVE: Software (lead), Investigation (lead), Visualization (supporting); IVV: Formal Analysis (lead), Writing – Review & Editing (supporting).
@@ Appendix A Detailed derivation of the master equation
Let the ODF of the conductor segment lengths be $f_{l}(l_{k};N_{\text{j}})$. Multiply equation (34) by $f_{l}(l_{k};N_{\text{j}})$ and integrate from 0 to $l$, then